# No-prior Bayes reIMagined: probabilistic approximations of possibilistic inferential models

### v2

@@ Abstract
When prior information is lacking, the go-to strategy for probabilistic inference is to combine a “default prior” and the likelihood via Bayes’s theorem. Objective Bayes, (generalized) fiducial inference, etc. fall under this umbrella. This construction is natural, but the corresponding posterior distributions generally only offer limited, approximately valid uncertainty quantification. The present paper takes a reimagined approach offering posterior distributions with stronger reliability properties. The proposed construction starts with an inferential model (IM), one that takes the mathematical form of a data-driven possibility measure and features exactly valid uncertainty quantification, and then returns a so-called inner probabilistic approximation thereof. This inner probabilistic approximation inherits many of the original IM’s desirable properties, including credible sets with exact coverage and asymptotic efficiency. The approximation also agrees with the familiar Bayes/fiducial solution obtained in applications where the model has a group transformation structure. A Monte Carlo method for evaluating the probabilistic approximation is presented, along with numerical illustrations.
@@ 1 Introduction
Fisher, (1933); Fisher, 1935a ; Fisher, 1935b and his fiducial argument was the first proposed solution to the above problem, making “a bold attempt to make the Bayesian omelet without breaking the Bayesian egg” (Savage, 1961); see Zabell, (1992) for an excellent survey. The consensus is that Fisher’s solution failed, but even his “biggest blunder” (Efron, 1998) had a major impact, inspiring fundamental advances such as confidence limits (Neyman, 1941), imprecise probability (Dempster, 1966), and new proposed solutions, including generalized fiducial inference (Hannig et al., 2016; Murph et al., 2021), confidence distributions (Xie and Singh, 2013; Thornton and Xie, 2020; Cox, 2006; Schweder and Hjort, 2016), and objective/non-informative Bayes (Jeffreys, 1946; Berger, 2006; Berger et al., 2009; Kass and Wasserman, 1996). In what follows, to keep terminology simple, I’ll (inaccurately) refer to all of these approaches as no-prior Bayes solutions.
@@
A difficulty with Bayesian-like probabilistic inference in this no-prior setting concerns the interpretation of the posterior probabilities themselves. When genuine prior information is available, then the Bayesian posterior probabilities are the unique coherent updates of prior beliefs given the observed data. But when there’s no prior beliefs to update, and a default prior is used in its place, of course the aforementioned “updating” interpretation disappears and, consequently, it’s unclear if the corresponding posterior distribution is meaningful at all. Indeed, “[Bayes’s formula] does not create real probabilities from hypothetical probabilities” (Fraser, 2014) and, put a bit more forcefully,
@@
…any serious mathematician would surely ask how you could use [Bayes’s theorem] with one premise missing by making up an ingredient and thinking that the conclusions of the [theorem] were still available. (Fraser, 2011b )
@@
Fortunately, coherent updating of prior beliefs isn’t necessary for a posterior to have belief-forming inferential weight. But making this case requires, first, defining what it takes for a framework to have belief-forming inferential weight and, second, demonstrating that a proposed framework meets this requirement. To me, the only way forward is to demonstrate that the posterior probabilities are reliable, i.e., it’s a demonstrably rare event that the posterior assigns high (resp. low) probabilities to assertions about the unknown that are false (resp. true). The advantage of such a requirement is that the belief-forming inferential weight is inherited from Fisher’s inductive logic: if, e.g., assigning high probability to a false assertion is a rare event and, in the application at hand, the assertion is assigned high probability, then it’s safe to infer that the assertion is true because rare events effectively don’t happen. No-prior Bayes solutions often have some form of reliability, but the false confidence theorem (Balch et al., 2019) says that none have the strong reliability property I just described; see, also, Martin, (2019); Martin, 2024b and Section 2 below. Since no-prior Bayes solutions can’t offer reliable belief assignments, there’s motivation to look beyond these existing approaches.
@@
An alternative to these probabilistic approaches is the inferential model (IM) framework (e.g., Martin and Liu, 2013; Martin and Liu, 2015a ; Martin, 2022b ; Martin, 2024a ), which returns a data-dependent possibility measure offering reliable uncertainty quantification about unknown parameters. Further details and references about possibility measures and the specific IM construction are given in Section 2 below. Importantly, my proposed shift from probabilistic to possibilistic uncertainty quantification facilitates strong reliability and, among other things, this strong reliability implies the usual frequentist error rate control without giving up on the fully conditional Bayesian-like reasoning.
@@
Despite the obvious benefits, I’m under no delusion that statisticians will abandon probabilistic inference in favor of possibilsitic IMs in the near future. But that doesn’t mean IMs have to wait for the distant future to make a contribution. The “reIMagined” approach I’m proposing here starts with the provably reliable possibilistic IM and then reads off and returns an inner probabilistic approximation as a novel no-prior Bayes solution. As I’ll demonstrate below, the proposed inner probabilistic approximation inherits some—but necessarily not all—of the original possibilistic IM’s strong reliability properties: in particular, the corresponding credible sets for the full parameter are exact confidence sets. There’s just only so much that can be achieved by constructing default priors and checking if the corresponding posterior distributions are reliable. This motivates a new approach, and my suggestion is to prioritize the reliability properties and directly construct a data-dependent probability that has those desired properties.
@@
The remainder of the paper is organized as follows. Some brief background on possibility theory and possibilistic IMs is presented in Section 2. A characterization of the possibilistic IM’s credal set is presented in Section 3, which is then used to define the corresponding inner probabilistic approximation. Various properties of the proposed inner probabilistic approximation are developed in Section 4, including a correspondence with the familiar no-prior Bayes solutions in models that feature a group-theoretic invariance property and a version of the celebrated Bernstein–von Mises theorem that establishes its asymptotic efficiency. A strategy for (approximately) computing the IM’s inner probabilistic approximation is presented in Section 5 and a new, valid, and efficient solution to the technically challenging and practically relevant Behrens–Fisher problem is presented in Section 6. Some concluding remarks are given in Section 7 and the appendices contain some additional technical details and illustrations.
@@ 2.1 Possibility theory
Possibility measures (e.g., Dubois and Prade, 1988) are among the simplest forms of imprecise probability, closely linked to fuzzy set theory (e.g., Zadeh, 1978) and Dempster–Shafer theory (e.g., Shafer, 1976, 1987). Some applications in statistics are described in Dubois, (2006), and below I’ll mention one possibility-theoretic approach to statistical inference that has a long history. Probability and possibility theory differ philosophically and mathematically, but here I’ll only briefly discuss the latter; for a discussion of the former, the presentation in Shackle, (1961) is clear and compelling.
@@
The mathematical differences between probability and possibility theory can be succinctly summarized as follows: optimization is to possibility theory what integration is to probability theory. That is, a possibility measure $\overline{\Pi}$ defined on a space $\mathbb{Z}$ is determined by a function $\pi:\mathbb{Z}\to[0,1]$ with the property that $\sup_{z\in\mathbb{Z}}\pi(z)=1$. This function is called the possibility contour and the supremum-equals-1 property is a normalization condition analogous to the familiar integral-equals-1 property of probability densities. Then the possibility measure is determined by optimizing its contour, i.e., $\overline{\Pi}(A)=\sup_{z\in A}\pi(z)$, for any $A\subseteq\mathbb{Z}$, just like a probability measure is determined by integrating its density.
@@
The different calculus has a number of implications. Of particular relevance for the developments in this paper is that the aforementioned supremum-equals-1 normalization condition ensures that $\overline{\Pi}$ is a coherent upper probability (e.g., De Cooman, 1997; De Cooman and Aeyels, 1999; Walley, 1997) in the spirit of Walley, (1991) and others. Among other things, this means that $\overline{\Pi}$ determines a non-empty (closed and convex) set of ordinary probabilities that it dominates:
$\mathscr{C}(\overline{\Pi})=\{\mathsf{Q}\in\text{probs}(\mathbb{Z}):\mathsf{Q}(H)\leq\overline{\Pi}(H)\text{ for all measurable $H$}\},$ (1)
@@
where $\text{probs}(\mathbb{Z})$ is the set of probabilities supported on (the Borel $\sigma$-algebra of measurable subsets of) $\mathbb{Z}$. The set $\mathscr{C}(\overline{\Pi})$ is called the credal set and all (coherent) upper probabilities have one. Aside from being relatively simple, a key advantage to possibilistic uncertainty quantification is that the associated credal set has a statistically oriented characterization, which is important for the statistical developments below.
@@ 2.2 Possibilistic IMs
Consider a model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ consisting of probability distributions supported on a sample space $\mathbb{X}$, with parameter space $\mathbb{T}$. Suppose that the observable data $X$, taking values in $\mathbb{X}$, is a sample from the distribution $\mathsf{P}_{\Theta}$, where $\Theta\in\mathbb{T}$ is the unknown/uncertain “true value.” Prior information about $\Theta$ is assumed to be vacuous. The model and observed data $X=x$ determine a relative likelihood function
$R(x,\theta)=\frac{p_{\theta}(x)}{\sup_{\vartheta}p_{\vartheta}(x)},\quad\theta\in\mathbb{T}.$
I will assume throughout that the denominator is finite for almost all $x$.
The relative likelihood defines a possibility contour, i.e., a non-negative function such that $\sup_{\theta}R(x,\theta)=1$ for almost all $x$. This contour determines a possibility measure that can be used for data-driven uncertainty quantification about $\Theta$, which has been extensively studied in (e.g., Shafer, 1982; Wasserman, 1990a ; Denœux, 2006, 2014). This likelihood-driven possibility has a number of desirable properties. What it lacks, however, is a justification for why the “possibilities” assigned to hypotheses about $\Theta$ have belief-forming inferential weight. Fortunately, it’s straightforward to achieve the desired goal by “validifying” (Martin, 2022a ; Martin, 2022b ) the relative likelihood; see, also, Martin, (2015, 2018). This amounts to applying a version of the probability-to-possibility transform (e.g., Dubois et al., 2004; Hose, 2022), and the result is the possibilistic IM’s contour function:
@@
The corresponding possibility measure is defined via optimization as
$\overline{\Pi}_{x}(H)=\sup_{\theta\in H}\pi_{x}(\theta),\quad H\subseteq\mathbb{T}.$ (3)
@@
Figure 1: Illustration of a typical possibility contour $\theta\mapsto\pi_{x}(\theta)$ and how the possibility measure $\overline{\Pi}_{x}$ is determined from it. Here, the hypothesis $H$ is the interval $[3,5]$ and the maximum value $\overline{\Pi}_{x}(H)$ on contour is highlighted. Horizontal dashed line at $\alpha=0.1$ determines the 90% confidence interval $C_{0.1}(x)$ in (5).
As in Section 1, an essential feature of this IM construction is its validity property:
@@
This has a number of important consequences. First, (4) immediately implies that the upper $\alpha$ level set of the possibility contour,
$C_{\alpha}(x)=\{\theta\in\mathbb{T}:\pi_{x}(\theta)\geq\alpha\},\quad\alpha\in[0,1],$ (5)
is a $100(1-\alpha)$confidence set, i.e., $\sup_{\theta\in\mathbb{T}}\mathsf{P}_{\theta}\bigl{\{}C_{\alpha}(X)\not\ni\theta\bigr{\}}\leq\alpha$. Figure 1 gives a depiction of $C_{\alpha}(x)$, with $\alpha=0.1$. Second, from (3) and (4),
@@
$\sup_{\theta\in H}\mathsf{P}_{\theta}\bigl{\{}\overline{\Pi}_{X}(H)\leq\alpha\bigr{\}}\leq\alpha,\quad\text{$\alpha\in[0,1]$, $H\subseteq\mathbb{T}$}.$ (6)
@@
In words, a valid IM assigns possibility $\leq\alpha$ to true hypotheses at rate $\leq\alpha$ as a function of data $Z$. This gives the IM its “inferential weight”—(6) implies that $\overline{\Pi}_{x}(H)$ is not expected to be small when $H$ is true, so one is inclined to doubt the truthfulness of a hypothesis $H$ if $\overline{\Pi}_{x}(H)$ is small, and the strength of this inclination is determined by the magnitude of $\overline{\Pi}_{x}(H)$. This of course can be turned into a formal test procedure with frequentist error rate guarantees: by (6), the test “reject $H$ if $\overline{\Pi}_{x}(H)\leq\alpha$” controls the Type I error probability at level $\alpha$. Third, the above property ensures that the possibilistic IM is safe from false confidence (Balch et al., 2019; Martin, 2019; Martin, 2024b ), unlike default-prior Bayes and fiducial solutions. Further insights about IMs are presented in Appendix A.
@@
The IM output is a coherent imprecise probability and, consequently, there is an associated (non-empty) credal set, $\mathscr{C}(\overline{\Pi}_{x})$ as in (1). The members of $\mathscr{C}(\overline{\Pi}_{x})$, which I’ll denote by $\mathsf{Q}_{x}$, depend on $x$ because $\overline{\Pi}_{x}$ does, but need not correspond to Bayesian posterior distributions under any prior. Fortunately, an interpretation can be given to the members of $\mathscr{C}(\overline{\Pi}_{x})$, and this is key to my developments in Section 3.

### v3

@@ Abstract
When prior information is lacking, the go-to strategy for probabilistic inference is to combine a “default prior” and the likelihood via Bayes’s theorem. Objective Bayes, (generalized) fiducial inference, etc. fall under this umbrella. This construction is natural, but the corresponding posterior distributions generally only offer limited, approximately valid uncertainty quantification. The present paper takes a reimagined approach that yields posterior distributions with stronger reliability properties. The proposed construction starts with an inferential model (IM), one that takes the mathematical form of a data-driven possibility measure and features exactly valid uncertainty quantification, and then returns a so-called inner probabilistic approximation thereof. This inner probabilistic approximation inherits many of the original IM’s desirable properties, including credible sets with exact coverage and asymptotic efficiency. The approximation also agrees with the familiar Bayes/fiducial solution in applications where the model has a group invariance structure. A Monte Carlo method for evaluating the probabilistic approximation is presented, along with numerical illustrations.
@@ 1 Introduction
Fisher, (1933); Fisher, 1935a ; Fisher, 1935b and his fiducial argument was the first proposed solution to the above problem, “a bold attempt to make the Bayesian omelet without breaking the Bayesian egg” (Savage, 1961); see Zabell, (1992) and Dawid, (2024) for excellent surveys. The consensus is that Fisher’s solution failed, but even his “biggest blunder” (Efron, 1998) had a major impact, inspiring fundamental advances such as confidence limits (Neyman, 1941), imprecise probability (Dempster, 1966), and new proposed solutions, including generalized fiducial inference (Hannig et al., 2016; Murph et al., 2024), confidence distributions (Xie and Singh, 2013; Thornton and Xie, 2020; Cox, 2006; Schweder and Hjort, 2016), and objective/non-informative Bayes (Jeffreys, 1946; Berger, 2006; Berger et al., 2024). In what follows, to keep terminology simple, I’ll refer to all of these approaches as no-prior Bayes solutions.
@@
A difficulty with Bayesian-like probabilistic inference in this no-prior setting concerns the interpretation of the posterior probabilities themselves. When genuine prior information is available, then the Bayesian posterior probabilities are the unique coherent updates of prior beliefs given the observed data. But when there’s no prior beliefs to update, and a default prior is used in its place, of course the aforementioned “updating” interpretation disappears and, consequently, it’s unclear if the corresponding posterior distribution is meaningful at all. Indeed, “[Bayes’s formula] does not create real probabilities from hypothetical probabilities” (Fraser, 2014) and, more forcefully,
@@
…any serious mathematician would surely ask how you could use [Bayes’s theorem] with one premise missing by making up an ingredient and thinking that the conclusions of the [theorem] were still available (Fraser, 2011b ).
@@
Fortunately, coherent updating of prior beliefs isn’t necessary for a posterior to have belief-forming inferential weight. But making this case requires, first, defining what it takes for a framework to have belief-forming inferential weight and, second, demonstrating that a proposed framework meets this requirement. To me, the only way forward is to demonstrate that the posterior probabilities are reliable, i.e., it’s a demonstrably rare event that the posterior assigns high (resp. low) probabilities to assertions about the unknown that are false (resp. true). The advantage of such a requirement is that the belief-forming inferential weight is inherited from Fisher’s inductive logic: if, e.g., assigning low probability to a true assertion is a rare event and, in the application at hand, the assertion is assigned low probability, then it’s safe to infer that the assertion is false because rare events effectively don’t happen. No-prior Bayes solutions often have some form of reliability, but the false confidence theorem (Balch et al., 2019; Martin, 2019) says that none have the strong reliability property just described. Since no-prior Bayes solutions can’t offer reliable belief assignments, there’s motivation to look beyond these existing approaches.
@@
Inferential models (IMs) (e.g., Martin and Liu, 2013, 2015; Martin, 2025b ) are an alternative to the aforementioned probabilistic approaches, whose output is a data-dependent possibility measure that offers provably reliable uncertainty quantification about unknown parameters. Further details about possibility measures and a specific IM construction are given in Section 2. Importantly, my proposed shift from probabilistic to possibilistic uncertainty quantification facilitates strong reliability and, among other things, this strong reliability implies the usual frequentist error rate control without giving up on fully conditional Bayesian-like reasoning.
@@
Despite the benefits of possibilistic IMs compared to probabilistic no-prior Bayes solutions, I’m under no delusion that statisticians will abandon probabilism in favor of possibilism in the near future. But that doesn’t mean IMs have to wait for the distant future to make a contribution. The “reIMagined” approach I’m proposing here starts with the provably reliable, likleihood-based possibilistic IM and then reads off and returns an inner probabilistic approximation as a novel no-prior Bayes solution. As I’ll demonstrate below, the proposed inner probabilistic approximation inherits some—but necessarily not all—of the original possibilistic IM’s strong reliability properties: in particular, the corresponding highest posterior density credible sets for the full parameter are confidence sets. The motivation for this new approach is that there’s only so much that can be achieved by constructing default priors and checking if the corresponding posterior distributions are reliable. My suggestion here is to prioritize the reliability properties and directly construct a data-dependent probability that has those desired properties.
@@
The remainder of the paper is organized as follows. Some background on possibility theory and IMs is presented in Section 2. A characterization of the IM’s credal set is presented in Section 3, which is then used to define the corresponding inner probabilistic approximation. Various properties of the proposed inner probabilistic approximation are developed in Section 4, including an agreement with the familiar no-prior Bayes solutions in invariant models and a version of the celebrated Bernstein–von Mises theorem that establishes its asymptotic efficiency. A strategy for (approximately) computing the IM’s inner probabilistic approximation is presented in Section 5 and a new, valid, and efficient solution to the technically challenging and practically relevant Behrens–Fisher problem is presented in Section 6. Some concluding remarks are given in Section 7 and the appendices contain some additional technical details and illustrations.
@@ 2.1 Possibility theory
Possibility measures (e.g., Dubois and Prade, 1988) are among the simplest imprecise probability models, closely linked to fuzzy set theory (e.g., Zadeh, 1978) and Dempster–Shafer theory (e.g., Shafer, 1976, 1987). Applications in statistics are described in Dubois, (2006) and Dubois and Denœux, (2010); see, also, Section 2.2.
@@
The mathematical differences between probability and possibility theory can be succinctly summarized as follows: optimization is to possibility theory what integration is to probability theory. That is, a possibility measure $\mathsf{\Pi}$ defined on a space $\mathbb{Z}$ is determined by a function $\pi:\mathbb{Z}\to[0,1]$ with the property that $\sup_{z\in\mathbb{Z}}\pi(z)=1$. This function is called the possibility contour and the supremum-equals-1 property is a normalization condition analogous to the familiar integral-equals-1 property of probability densities. Then the possibility measure is determined by optimizing its contour, i.e., $\mathsf{\Pi}(A)=\sup_{z\in A}\pi(z)$, for any $A\subseteq\mathbb{Z}$, just like a probability measure is determined by integrating its density.
@@
The different calculus has a number of implications. Of particular relevance for the developments in this paper is that the aforementioned supremum-equals-1 normalization condition ensures that $\mathsf{\Pi}$ is a coherent upper probability (e.g., De Cooman, 1997; De Cooman and Aeyels, 1999) in the spirit of Walley, (1991) and others. Among other things, this means that $\mathsf{\Pi}$ determines a non-empty (closed and convex) set of probabilities it dominates:
$\mathscr{C}(\mathsf{\Pi})=\{\mathsf{Q}\in\mathcal{P}(\mathbb{Z}):\mathsf{Q}\preceq\mathsf{\Pi}\},$ (1)
@@
where $\mathcal{P}(\mathbb{Z})$ is the set of probabilities supported on the Borel $\sigma$-algebra of measurable subsets of $\mathbb{Z}$ and $\mathsf{Q}\preceq\mathsf{\Pi}$” means $\mathsf{Q}(H)\leq\mathsf{\Pi}(H)$ for all measurable HH.” The set $\mathscr{C}(\mathsf{\Pi})$ is called the credal set corresponding to $\mathsf{\Pi}$ and all (coherent) upper probabilities have one. Aside from being relatively simple, a key advantage to possibilistic uncertainty quantification is that the associated credal set has a statistically oriented characterization, which is important for the statistical developments below.
Critical to the developments in the present paper, and closely tied to the above credal sets, is a notion of approximating a probability by a possibility and vice versa. I’ll focus here on the former, and leave the latter for Section 3. Suppose that $\mathsf{Q}$ is a probability on $\mathbb{T}$ with density/mass function $q$. The probability-to-possibility transform (e.g., Dubois et al., 2004; Hose, 2022) applied to $\mathsf{Q}$ yields a possibility measure $\mathsf{\Pi}$ determined by the contour
$\pi(\theta)=\mathsf{Q}\{q(\Theta)\leq q(\theta)\},\quad\theta\in\mathbb{T}.$
This $\mathsf{\Pi}$ has two key properties: first, $\mathscr{C}(\mathsf{\Pi})$ contains $\mathsf{Q}$, so that $\mathsf{\Pi}\succeq\mathsf{Q}$; second, $\mathsf{\Pi}$ is the “least imprecise” possibilistic approximation of $\mathsf{Q}$ in the sense that it has the smallest credal set among other suitable possibilistic approximations. More formally, $\mathscr{C}(\mathsf{\Pi})=\bigcap\mathscr{C}(\mathsf{\Omega})$, where the intersection is over all those possibility measures $\mathsf{\Omega}$ that satisfy $\mathsf{\Omega}\succeq\mathsf{Q}$ and whose associated contour $\omega$ is comonotone [^2] Comonotonicity here means that the contour $\omega(\theta)$ is an increasing function of $q(\theta)$. This constraint is motivated by a basic and fundamental principle in this possibility-theory context: “what is probable must be plausible” (Dubois and Prade, 1988, p. 121). with $q$. So, $\mathsf{\Pi}$ is often called an outer possibilistic approximation of $\mathsf{Q}$ (e.g., Dubois and Prade, 1990).
@@ 2.2 Possibilistic IMs
An inferential model (IM) is a mapping from data, model, etc. to an imprecise-probabilistic quantification of uncertainty about relevant unknowns. Critical to this approach and perspective is that the IM’s uncertainty quantification be reliable, as a function of data, in a sense to be described below. It’s this insistence on reliability that implies the IM output must take on the form of an imprecise probability. The first IM developments were built up using random sets and described using belief function terminology. More recent developments directly appeal to possibility-theoretic tools and reasoning in the IM construction and interpretation, and I refer to these as possibilistic IMs; see Martin, 2025b for more details. This paper focuses on this latter brand of IMs.
Consider a model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ consisting of probability distributions supported on a sample space $\mathbb{X}$, with parameter space $\mathbb{T}$. Suppose that the observable data $X$, taking values in $\mathbb{X}$, is a sample from the distribution $\mathsf{P}_{\Theta}$, where $\Theta\in\mathbb{T}$ is the unknown/uncertain “true value.” Prior information about $\Theta$ is assumed to be vacuous, but see Martin, 2022b for generalizations. The model and observed data $X=x$ determine a relative likelihood function
$R(x,\theta)=\frac{p_{\theta}(x)}{\sup_{\vartheta}p_{\vartheta}(x)},\quad\theta\in\mathbb{T},$
where $p_{\theta}$ is the density of $\mathsf{P}_{\theta}$. I will assume throughout that the denominator is finite for almost all $x$.
The relative likelihood itself defines a possibility contour, i.e., a non-negative function such that $\sup_{\theta}R(x,\theta)=1$ for almost all $x$, which can be used for data-driven uncertainty quantification about $\Theta$. This has been extensively studied in (e.g., Shafer, 1982; Wasserman, 1990a ; Denœux, 2006, 2014), and it has a number of desirable properties. What it lacks, however, is a formal calibration property justifying that the “possibilities” assigned to hypotheses about $\Theta$ have belief-forming inferential weight.
Fortunately, this reliability-focused calibration is easy to achieve, by “validifying” (Martin, 2022a ; Martin, 2022b ) the relative likelihood. That is, the likelihood-based possibilistic IM construction proceeds by applying a version of the probability-to-possibility transform to the relative likelihood, returning contour function:
@@
The corresponding possibility measure $\mathsf{\Pi}_{x}$ is defined via optimization just like I explained above:
$\mathsf{\Pi}_{x}(H)=\sup_{\theta\in H}\pi_{x}(\theta),\quad H\subseteq\mathbb{T}.$ (3)
@@
Figure 1: Illustration of a typical possibility contour $\theta\mapsto\pi_{x}(\theta)$ and how the possibility measure $\mathsf{\Pi}_{x}$ is determined from it. Here, the hypothesis $H$ is the interval $[3,5]$ and the maximum value $\mathsf{\Pi}_{x}(H)$ on contour is highlighted. Horizontal dashed line at $\alpha=0.1$ determines the 90% confidence interval $C_{0.1}(x)$ in (5).
As in Section 1, an essential feature of this possibilistic IM is its validity property:
@@
In words, (4) means that it’s a rare event for the IM to assign small possibility $\pi_{X}(\Theta)$ to the true $\Theta$, which, of course, was important to the inductive logic explained in Section 1. This has a number of important practical consequences as well. First, (4) immediately implies that the upper $\alpha$ level set of the possibility contour,
$C_{\alpha}(x)=\{\theta\in\mathbb{T}:\pi_{x}(\theta)>\alpha\},\quad\alpha\in[0,1],$ (5)
is a $100(1-\alpha)$% confidence set, i.e.,
$\sup_{\theta\in\mathbb{T}}\mathsf{P}_{\theta}\bigl\{C_{\alpha}(X)\not\ni\theta\bigr\}\leq\alpha.$
Figure 1 gives a depiction of $C_{\alpha}(x)$, with $\alpha=0.1$. Second, from (3) and (4),
@@
$\sup_{\theta\in H}\mathsf{P}_{\theta}\bigl\{\mathsf{\Pi}_{X}(H)\leq\alpha\bigr\}\leq\alpha,\quad\text{$\alpha\in[0,1]$, $H\subseteq\mathbb{T}$}.$ (6)
@@
In words, a valid IM assigns possibility $\leq\alpha$ to true hypotheses at rate $\leq\alpha$ as a function of data $Z$. This gives the IM its “inferential weight”—(6) implies that $\mathsf{\Pi}_{x}(H)$ is not expected to be small when $H$ is true, so one is inclined to doubt the truthfulness of a hypothesis $H$ if $\mathsf{\Pi}_{x}(H)$ is small, and the strength of this inclination is determined by the magnitude of $\mathsf{\Pi}_{x}(H)$. This of course can be turned into a formal test procedure with frequentist error rate guarantees: by (6), the test “reject $H$ if $\mathsf{\Pi}_{x}(H)\leq\alpha$” controls the Type I error probability at level $\alpha$. Third, the above property ensures that the possibilistic IM is safe from false confidence (Balch et al., 2019; Martin, 2019), unlike other no-prior Bayes solutions. Further insights about IMs are presented in Appendix A in the supplementary materials; see, also, Martin, 2025b .
@@
The IM output is a coherent imprecise probability and, consequently, there is an associated (non-empty) credal set, $\mathscr{C}(\mathsf{\Pi}_{x})$ as in (1). The members of $\mathscr{C}(\mathsf{\Pi}_{x})$, which I’ll denote by $\mathsf{Q}_{x}$, need not correspond to Bayesian posterior distributions under any prior. Fortunately, an interpretation can be given to the members of $\mathscr{C}(\mathsf{\Pi}_{x})$, and this is key to my developments in Section 3.
I’ll end this background section with three technical remarks about the IM scope and contribution. First, the reader certainly recognizes the contour function $\pi_{x}(\theta)$ in (2) as a p-value function corresponding to the likelihood ratio statistic. Aside from p-value function, this has appeared in the literature under many other names, e.g., preference functions (Spjøtvoll, 1983), significance functions (Fraser, 1991), and confidence curves (Blaker and Spjøtvoll, 2000; Birnbaum, 1961; Schweder and Hjort, 2002, 2016; Xie and Singh, 2013). A key point distinguishing IMs from this literature is that it embraces the inherent imprecise and leverages relevant imprecise-probabilistic results. That is, an IM determines a coherent and fully conditional imprecise-probabilistic quantification of uncertainty for inference and more (Martin, 2025b ). Even the more Bayesian-leaning developments along these lines (e.g., Pereira and Stern, 2022; Grünwald, 2023; Cortinovis and Caron, 2024) don’t take full advantage of what imprecise probability has to offer.
Second, with the big-picture assessment in the previous paragraph, one might compare IMs to other imprecise-probabilistic solutions like that in Walley, (1991). A critical observation is that, when prior information about $\Theta$ is vacuous, as assumed here, Walley’s generalized Bayes posterior is also vacuous, i.e., no learning is achieved. Since Bayesian reasoning isn’t fully satisfactory in such cases, new ideas, like those reviewed above, are needed to achieve inference that’s both valid and efficient.
Finally, with the connection to p-values above, it should be clear that other rankings besides the relative likelihood can be used in the construction. My reason for focusing on the relative likelihood here is two-fold: first, this is a principled choice in the case where the full parameter $\Theta$ is of interest (Martin, 2022b , Sec. 4); second, the focus of the present paper is on connections to Bayesian inference, and some direct dependence on the model’s likelihood function is surely needed to make that connection. But there are cases when it’s prudent to replace the relative likelihood with something else, e.g., when nuisance parameters are present as in Section 6.